1/2+(2/3 / 3/4)-(4/5 * 5/6)
Answer by lotanez(23)

1/2 + (2/3 / 3/4) - (4/5 * 5/6)
(do the division and multiplication first)
1/2 + (2/3 * 4/3) - (4/5 * 5/6)
1/2 + 8/9 - 20/30
(reduce 20/30 by dividing by 10 = 2/3)
1/2 + 8/9 - 2/3
(the LCD is 18)
(9)1/(9)2 = 9/18
(2)8/(2)9 = 16/18
(6)-2/(6)3 = -12/18
9/18 + 16/18 - 12/18 = 13/18